Suppose GRE Verbal scores are normally distributed with a mean of 459 and a standard deviation of 120. A university plans to offer tutoring jobs to students whose scores are in the top 14%. What is the minimum score required for the job offer?
Found 2 solutions by stanbon, mathmate:
Answer by stanbon(75887)

Suppose GRE Verbal scores are normally distributed with a mean of 459 and a standard deviation of 120. A university plans to offer tutoring jobs to students whose scores are in the top 14%. What is the minimum score required for the job offer?
-----
Find the z-score with a left tail of 0.86
invNorm(0.86) = 1.0803
-------
Solve for "x" using x = z*s + u
-----
x = 1.0803*120 + 459 = 588.64
=================
Cheers,
Stan H.
------------------

Answer by mathmate(429)


Question:
Suppose GRE Verbal scores are normally distributed with a mean of 459 and a standard deviation of 120. A university plans to offer tutoring jobs to students whose scores are in the top 14%. What is the minimum score required for the job offer?

Solution:
Top 14% means that 86% of the students are below.
The corresponding Z-score can be read from the Normal Distribution table as:
P(Z≤1.0803)=86% (left-tail only).
So convert to number of students,
Z=(X-μ)/σ
means that
X=Z*σ+&mu=1.0803*120+459=588.6
So minimum score is 588.6 to be in the top 14%.